OpenStudy (nastech):

Please help me solve this differential equation yI=(1-x)y2 + (2x-1) y-1 given y͵=1

OpenStudy (anonymous):

lol im learning this right now...

OpenStudy (turingtest):

\[y'=(1-x)y^2+(2x-1)y-1\text{ given }y=1\]correct? are you doing Laplace right now?

OpenStudy (turingtest):

\[y'=(1-x)y^2+(2x-1)y-1\text{ given }y(0)=1\]and what do you mean y=1, do you mean this^

OpenStudy (nastech):

no, Bernoulli and Riccati's equations

OpenStudy (turingtest):

I don't know Riccati and I have to review Bernoulli...

OpenStudy (turingtest):

do I have the problem right?\[y'=(1-x)y^2+(2x-1)y-1\text{ given }y(0)=1\]

OpenStudy (turingtest):

or is it\[y'=(1-x)y^2+(2x-1)y^{-1}\text{ given }y(0)=1\]

OpenStudy (anonymous):

divide the original equation by highest power of Y, example if we have a y^2 term we divide it by Y^2 then rewrite it, let 1/y= K OR variable of your choice sub that in the equation and that will give you a hint how to proceed further

OpenStudy (anonymous):

must say that if you have chosen 1/y= k, then make sure you find k',